Question One

Matrix P = ( 3 -1 ) and Q = ( 3 4 ) If P and Q have equal determinants find the value of x.
(-2 0 ) (-4 x )

Question Two

Factorise Completely 5X^2 - X - 4
1
Expert's answer
2020-04-08T12:41:08-0400

Solution (1)

Determinant of matrix P

"det \space \begin{pmatrix}3&-1\\ -2&0\end{pmatrix}=0-(-1)(-2)=-2"

Determinant of matrix Q

"det \:\begin{pmatrix}3&4\\ \:\:-4&x\end{pmatrix}=3(x)-4(-4)=3x+16"

From question |P| = |Q|

"\Rightarrow -2=3x+16"

"\Rightarrow 3x=-18"

"\Rightarrow x=-18/3 = -6"


Solution (2)

"5x^2-x-4"

"=5x^2-5x+4x-4"

"=5x(x-1)+4(x-1)"

"=(5x+4)(x-1)"